# FILP-3D: Enhancing 3D Few-shot Class-incremental Learning with Pre-trained Vision-Language Models

### v1

@@ Abstract
Few-shot class-incremental learning (FSCIL) aims to mitigate the catastrophic forgetting issue when a model is incrementally trained on limited data. While the Contrastive Vision-Language Pre-Training (CLIP) model has been effective in addressing 2D few/zero-shot learning tasks, its direct application to 3D FSCIL faces limitations. These limitations arise from feature space misalignment and significant noise in real-world scanned 3D data. To address these challenges, we introduce two novel components: the Redundant Feature Eliminator (RFE) and the Spatial Noise Compensator (SNC). RFE aligns the feature spaces of input point clouds and their embeddings by performing a unique dimensionality reduction on the feature space of pre-trained models (PTMs), effectively eliminating redundant information without compromising semantic integrity. On the other hand, SNC is a graph-based 3D model designed to capture robust geometric information within point clouds, thereby augmenting the knowledge lost due to projection, particularly when processing real-world scanned data. Considering the imbalance in existing 3D datasets, we also propose new evaluation metrics that offer a more nuanced assessment of a 3D FSCIL model. Traditional accuracy metrics are proved to be biased; thus, our metrics focus on the model’s proficiency in learning new classes while maintaining the balance between old and new classes. Experimental results on both established 3D FSCIL benchmarks and our dataset demonstrate that our approach significantly outperforms existing state-of-the-art methods. Code is available at Code is available at https://github.com/HIT-leaderone/FLIP-3D
@@ 1 Introduction
The proliferation of 3D content in recent years has been marked by both synthetic creations [2, 5] and real-world reconstruction [29, 24, 35]. This expansion in data scale naturally introduces new classes into the 3D domain. However, these emerging categories often comprise a limited number of instances, posing a challenge for existing 3D recognition models. Consequently, 3D few-shot class-incremental learning (FSCIL) has become increasingly important in practical applications.
Previous work in 2D few-shot and zero-shot tasks [20, 47] has shown that Pre-Training Models (PTMs) excel in incremental learning scenarios, often outperforming non-PTM-based approaches. This superior performance is largely attributed to the prior knowledge that PTMs acquire [6, 10, 21], which enhances generalization in downstream tasks. Motivated by these successes, we aim to leverage PTMs to imbue our model with shape-related prior knowledge. However, the limited scale of available 3D data hampers the effectiveness of existing 3D PTMs in downstream tasks. To address this, recent studies [43, 11, 40] have successfully aligned 3D representations with Vision-Language (V-L) PTM knowledge. Building on this, we map point cloud data into a cross-modal space, utilizing vision and language pre-training to improve performance in 3D FSCIL tasks. Specifically, we generate multi-view depth maps from point clouds and employ template text prompts for classification, such as “an image of a {class name}”.
@@
While the integration of Vision-Language (V-L) Pre-Training Models (PTMs) into FSCIL tasks has yielded improvements over previous methods, two key challenges remain to limit performance. First, V-L PTMs are designed to capture detailed visual features such as color and texture. However, point cloud data, being a collection of discrete points, lacks the information to represent these intricate visual details. This discrepancy can lead to inaccurate feature extraction, adversely affecting performance. Second, the use of V-L PTMs amplifies the issue of noise sensitivity, particularly in real-world scanned data. Consequently, classification performance suffers significantly if noise disrupts the rendering of depth maps and object contours.
To tackle the aforementioned challenges, we introduce Few-shot class Incremental Learning tasks with Pre-training on 3D (FILP-3D), a framework that employs CLIP as its backbone and incorporates two innovative components: Redundant Feature Eliminator (RFE) and Spatial Noise Compensator (SNC), as is shown in Figure 1. RFE serves as a specialized dimensionality reduction technique, designed to remove redundant features while preserving semantic content. By precisely compressing dimensions within the V-L PTMs, RFE facilitates better alignment between the feature spaces of point clouds and their corresponding embeddings. On the other hand, SNC is a graph-based 3D model tailored to extract robust geometric information from point clouds. This additional layer of information enhances the geometric relationships between object components, mitigating the loss of knowledge that occurs during projection, especially in noisy, real-world scanned data. As a result, our FILP-3D framework gains improved resilience against noise interference.
Furthermore, we observe that the numbers of samples for different classes in test datasets can vary by up to two orders of magnitude, leading to a significant imbalance. This imbalance skews traditional accuracy metrics, making them less reliable for comprehensive evaluation. To address this, we introduce new evaluation metrics, namely $\rm NCAcc$ and $\rm F_{FSCIL}$. These metrics are designed to assess the model’s ability to learn new classes effectively while maintaining a balance between old and new classes, thereby providing a more nuanced evaluation of the model performance.
@@
  • We pioneer the application of Vision-Language Pre-Training Models (V-L PTMs) to 3D Few-Shot Class-Incremental Learning (FSCIL) tasks, achieving performance gains over existing models. The general embedding provided by V-L PTM, along with its embedded prior knowledge, can complement the missing information in few-shot tasks and alleviate catastrophic forgetting during the incremental learning process.
@@ 2 Related Work
In recent years, many works have been proposed to classify 3D point cloud objects. PointNet [18] and PointNet++ [19] design the architecture to maintain natural invariances of the data. DGCNN [34] connects the point set into a graph and designs a local neighbor aggregation strategy.
Some works employ masked point modeling [7, 12] as a 3D self-supervised learning strategy to achieve great success. For example, Point BERT [39] uses a pre-trained tokenizer to predict discrete point labels, while Point MAE [17] and Point-M2AE [42] apply masked autoencoders to directly reconstruct the masked 3D coordinates.
[…]
@@ 2.2 Few-Shot Class-Incremental Learning
The FSCIL problem was proposed by Tao et al. [28]. Concretely, FSCIL aims at learning from severely insufficient samples incrementally while preserving already learned knowledge. TOPIC [28] uses a neural gas network to learn and preserve the topology of features. Subsequently, CEC [41] utilizes a graph model to propagate context information between classifiers for adaptation. Also, some models like FACT [46] try to use virtual prototypes to reserve for new ones, thus ensuring incremental learning ability. The SOTA method BiDist [45] utilizes a novel distillation structure to alleviate the effects of forgetting.
Recently, with the breakthroughs in 2D PTMs, a number of works attempt to leverage the vast knowledge acquired by 2D PTMs, which is effective in learning new concepts and alleviating the problem of forgetting [22, 47, 33]. With almost perfect performance, these models have attracted a lot of interest and attention. The preceding methods are all 2D methods, except for Chowdhury et al. [3]’s work which explores the FSCIL task in 3D point cloud data. However, Chowdhury et al. [3]’s work neither supplements knowledge for few-shot data nor addresses the high-noise nature of real-world scanned data, resulting in limited performance in 3D FSCIL tasks. In contrast, our FILP-3D successfully addresses the shortcomings mentioned above by introducing V-L PTMs and two newly proposed modules.
@@ 3.1 Problem Formulation
Assuming a sequence of $B$ tasks $\mathcal{D}=\{\mathcal{D}^{1},\mathcal{D}^{2},\dots,\mathcal{D}^{B}\}$, FSCIL methods incrementally recognize novel classes with a small amount of training data. For the $b$-th task $\mathcal{D}^{b}=\{(\mathbf{x}_{i}^{b},y_{i}^{b})\}_{i=1}^{n_{b}}$, we have $n_{b}$ training samples, in which an instance $\mathbf{x}_{i}^{b}$ has a class label $y_{i}^{b}$ and $y_{i}^{b}$ belongs to the label set $Y_{b}$. We stipulate that $Y_{b}\cap Y_{b^{\prime}}=\emptyset$ when $b\not=b^{\prime}$. After the $b$-th training task, trained models are required to classify test sets of all the previous tasks $\{\mathcal{D}^{1},\dots,\mathcal{D}^{b}\}$. Note that, in the 3D FSCIL setting, $\mathbf{x}_{i}^{b}\in\mathbb{R}^{P\times 3}$ denotes a point cloud object, where $P$ donates the number of points in a point cloud object. $\mathcal{D}^{1}$ is the base task with a large-scale 3D training dataset, while much fewer samples are included in the incremental task $\mathcal{D}^{b}$, i.e., $n_{b}\ll n_{1}$ for $b>1$.
@@ 3.2 Model Overview
Figure 2: Overview of FILP-3D. FILP-3D mainly consists of three components, i.e., 3D branch (SNC), 2D branch, and textual prototypes. Then, the 3D feature and the 2D feature will be fused as a global feature, and be used to calculate probability through the REF module alongside textual prototypes.
The framework of FILP-3D is shown in Figure 2. For each training sample $(\mathbf{x}_{i}^{b},y_{i}^{b})$, we mainly have three branches: 1) point cloud $\mathbf{x}_{i}^{b}$ is fed into 3D encoder to obtain an original point feature $\mathbf{f}^{3D}$ and then aligned to $\mathbf{f}^{p}$, 2) $\mathbf{x}_{i}^{b}$ is then rendered as multi-view depth maps ${\rm\mathbf{D}}^{b}_{i}$, embedded as depth features ${\rm\mathbf{F}}^{2D}$, and finally merged into a global depth feature $\mathbf{f}^{d}$, 3) visible class names are templated into text prompts and encoded as text features ${\rm\mathbf{F}}^{t}$. The point features $\mathbf{f}^{p}$ and the global depth feature $\mathbf{f}^{d}$ are then fused as a global feature $\mathbf{f}^{g}$. Afterward, we use pre-processed principal components $\mathbf{V}$ (refer to Sec. 5 in Suppl. for more details of pre-processing) to eliminate redundant dimensions of $\mathbf{f}^{g}$ and $\mathbf{F}^{t}$ and generate $\widetilde{\mathbf{f}}^{g}$ and $\widetilde{\mathbf{F}}^{t}$. The final predicted probability is calculated by our proposed renormalized cosine similarity.

### v2

@@ Abstract
Few-shot class-incremental learning (FSCIL) aims to mitigate the catastrophic forgetting issue when a model is incrementally trained on limited data. However, many of these works lack effective exploration of prior knowledge, rendering them unable to effectively address the domain gap issue in the context of 3D FSCIL, thereby leading to catastrophic forgetting. The Contrastive Vision-Language Pre-Training (CLIP) model serves as a highly suitable backbone for addressing the challenges of 3D FSCIL due to its abundant shape-related prior knowledge. Unfortunately, its direct application to 3D FSCIL still faces the incompatibility between 3D data representation and the 2D features, primarily manifested as feature space misalignment and significant noise. To address the above challenges, we introduce the FILP-3D framework with two novel components: the Redundant Feature Eliminator (RFE) for feature space misalignment and the Spatial Noise Compensator (SNC) for significant noise. RFE aligns the feature spaces of input point clouds and their embeddings by performing a unique dimensionality reduction on the feature space of pre-trained models (PTMs), effectively eliminating redundant information without compromising semantic integrity. On the other hand, SNC is a graph-based 3D model designed to capture robust geometric information within point clouds, thereby augmenting the knowledge lost due to projection, particularly when processing real-world scanned data. Moreover, traditional accuracy metrics are proven to be biased due to the imbalance in existing 3D datasets. Therefore we propose 3D FSCIL benchmark FSCIL3D-XL and novel evaluation metrics that offer a more nuanced assessment of a 3D FSCIL model. Experimental results on both established and our proposed benchmarks demonstrate that our approach significantly outperforms existing state-of-the-art methods. Code is available at: https://github.com/HIT-leaderone/FILP-3D
keywords:
3D classification, FSCIL, 3D from multi-view, V-L pre-trained model
journal: Pattern Recognition
[…]
[label1]addressline=Harbin Institute of Technology, city=Harbin, country=China
[…]
[label2]organization=Independent Researcher
@@ 1 Introduction
The advancement of deep learning has led to significant progress in class recognition within predefined label sets, as demonstrated by various groundbreaking models [1, 2, 3, 4]. However, in real-world settings, it’s common to encounter new types of data in a continuous stream, introducing classes that were not part of the original training. This situation is especially challenging in areas where collecting data is expensive, such as identifying rare animals [5]. In response, the research community has introduced several Few-Shot Class-Incremental Learning (FSCIL) models [6, 7, 8]. These models aim to gradually learn from a small number of examples of new concepts while keeping the knowledge previously gained. Although much of the existing research  [6, 7] has focused on image data, these approaches often fall short for FSCIL tasks in point cloud data. For example, a robot that has been trained extensively on synthetic data needs to adjust to new classes based on real-world point clouds. In this scenario, the domain gap between synthetic training data and real-world data makes feature-prototype alignments cumbersome during the incremental steps, eventually magnifying forgetting and overfitting issues [8]. Therefore, there is a critical need for 3D FSCIL models that not only tackle the issues of catastrophic forgetting and overfitting but also adapt to the domain gap encountered during incremental learning.
The pioneering work by  Chowdhury et al. [8] marks the first significant exploration into 3D FSCIL, offering a novel solution to this issue. They propose extracting principal components from base task instances, termed microshapes, which represent distinct aspects of 3D shapes. This method allows for the effective alignment of 3D point cloud features with the language prototypes of classes, facilitating more accurate class recognition. However, the microshapes primarily draw from knowledge of base task instances and lack supplementation of sufficient shape-related prior knowledge. Therefore, the features exacted by their model lack sufficient generalization capability. Moreover, the point features extracted by PointNet [2] are not in alignment with the textual prototypes. Thus,  [8] finds it challenging to simultaneously retain features associated with base classes (old knowledge) and integrate features of novel classes (new knowledge). In contrast, several methods [9, 10] have shown that the extensive prior knowledge embedded in Vision-Language (V-L) Pre-trained Models (PTMs) can significantly bridge the domain gap in domain adaptation tasks. Furthermore, Radford et al. [11] demonstrate that the general and adaptable embeddings provided by V-L PTMs are effective in addressing the dual challenges of forgetting and overfitting commonly encountered in FSCIL tasks. Given above insights, V-L PTMs emerge as highly promising backbones for overcoming the specific challenges associated with 3D FSCIL. Therefore, we attempt to introduce PTMs into the 3D FSCIL task, to address the lack of prior knowledge and the issue of naturally aligned encoders in previous 3D FSCIL works [8].
@@
The incorporation of V-L PTMs into Few-Shot Class-Incremental Learning (FSCIL) tasks, through direct projection techniques, has marked a significant advancement over previous methodologies according to our experiment (please refer to Section. 6 for more details). Despite these improvements, the incompatibility between 3D data representation and the 2D feature encoder impede further performance enhancements, primarily manifested as two aspects. First, point cloud data, which consists of discrete points in 3D space, primarily represents geometric features of objects. It inherently misses richer visual nuances, such as color and texture. This deficiency is especially significant given CLIP’s visual encoder, which inherently seeks and incorporates these visual details into its semantic representation [12]. Consequently, this feature space misalignment could harm the precise classification of depth maps. Second, point clouds derived from real-world scans often suffer from significant noise [13], which distorts the extraction of depth-centric features. This problem is accentuated in multi-view projection techniques, where a significant fraction of the points might be occluded, thereby increasing the potential noise [14]. Such disturbances can enlarge the domain gap and upset the balance in few-shot incremental learning, leading to overfitting [8].
To tackle the mismatch between 3D representation and V-L PTM’s 2D encoder in direct projection techniques [8, 15], we introduce Few-shot class Incremental Learning tasks with Pre-training on 3D (FILP-3D), a framework that employs CLIP as its backbone and incorporates two innovative components: Redundant Feature Eliminator (RFE) to tackle the feature space misalignment and Spatial Noise Compensator (SNC) to tackle the significant noise issue, as is shown in Figure 1. RFE serves as a specialized dimensionality reduction technique, designed to remove redundant features while preserving semantic content. By precisely compressing dimensions within the V-L PTMs, RFE facilitates better alignment between the feature spaces of point clouds and their corresponding embeddings. On the other hand, SNC is a graph-based 3D model tailored to extract robust geometric information from point clouds. This additional layer of information enhances the geometric relationships between object components, mitigating the loss of knowledge that occurs during projection, especially in noisy, real-world scanned data. As a result, our FILP-3D framework gains improved resilience against noise interference. Furthermore, we observe that the numbers of samples for different classes in test datasets can vary by up to two orders of magnitude, leading to a significant imbalance. This imbalance skews traditional accuracy metrics, making them less reliable for comprehensive evaluation. To address this, we introduce new evaluation metrics, namely $\rm NCAcc$ and $\rm F_{FSCIL}$. These metrics are designed to assess the model’s ability to learn new classes effectively while maintaining a balance between old and new classes, thereby providing a more nuanced evaluation of the model’s performance.
@@
  • 1.
    We pioneer the application of Vision-Language Pre-Training Models (V-L PTMs) to 3D Few-Shot Class-Incremental Learning (FSCIL) tasks, achieving performance gains over existing models. The general embedding provided by V-L PTM, along with its embedded prior knowledge, can complement the missing information in few-shot tasks, alleviate catastrophic forgetting, and narrow the domain gap during the incremental learning process.
  • 2.
@@ 2 Related Work
2.1 Vision-language Pre-training
Vision-language (V-L) pre-training has become a popular topic in multi-modal tasks. Pre-trained by large-scale image-text pairs, those models are capable of many downstream tasks, e.g., visual question answering, and text-to-image generation. CLIP [12] further leverages V-L pre-training to transfer cross-modal knowledge from the massive image-text pairs (400M), allowing the natural language to understand visual concepts, and exhibit strong capability in many tasks.  Zhu et al. [9], Kang et al. [10] utilize CLIP to reduce the domain gap via CLIP’s prior knowledge in the domain adaption task.  Radford et al. [11], Zhou et al. [16] leverage general and adaptive embeddings of CLIP to prevent forgetting during the incremental learning stage.  [15, 17, 3] verify that shape-related prior knowledge can enhance the model’s few-shot/zero-shot capability even in 3D point cloud modal.
Following CLIP, recent models like BLIP-2 [18] incorporated iterative learning strategies and efficient cross-attention mechanisms to refine vision-language representations further. More recent work focuses on expanding these models’ capabilities to handle video-language pre-training. For example, LLaVA [19] integrates large language models with visual inputs to enable multimodal dialogue and image understanding, marking a step forward in interactive VLM applications. These models collectively push the boundary of generalization, robustness, and adaptability in vision-language tasks, e.g., open-world video recognition [20].
@@ 2.1 Vision-language Pre-training
In recent years, many works have been proposed to classify 3D point cloud objects. PointNet [2] design the architecture to maintain natural invariances of the data. DGCNN [21] connects the point set into a graph and designs a local neighbor aggregation strategy.  Hao et al. [22] propose CGRL to enhance 3D point cloud classification by generating more discriminative features, particularly in zero-shot scenarios. And some works employ masked point modeling as a 3D self-supervised learning strategy to achieve great success. For example, Point BERT [23] uses a pre-trained tokenizer to predict discrete point labels, while Point-MAE [24] applies masked autoencoders to directly reconstruct the masked 3D coordinates.
Recently, inspired by the breakthroughs in V-L PTMs [12], a number of approaches have been suggested to transfer 2D PTM to point cloud tasks and show excellent performance. CLIP2Point [17] transfers CLIP to point cloud classification with image-depth pre-training. I2P-MAE [25] leverages knowledge of 2D PTMs to guide 3D MAE and ULIP [3] improves 3D understanding by aligning features from images, texts, and point clouds. These models are supported by prior knowledge from V-L PTMs, which can improve few-shot/zero-shot performance.
Although these models have achieved remarkable progress in the 3D few-shot/zero-shot domain, these 3D models still exhibit limited generalizability, e.g., they can hardly handle OOD examples like real-scanned data. Incremental learning can more effectively mine knowledge from real-world scenarios, thus benefiting downstream applications. In this work, we aim to introduce V-L PTMs to the 3D FSCIL task and further improve the generalizability and performance.
@@ 2.3 Few-Shot Class-Incremental Learning
The FSCIL problem was proposed by Tao et al. [6]. Concretely, FSCIL aims at learning from severely insufficient samples incrementally while preserving already learned knowledge. TOPIC [6] uses a neural gas network to learn and preserve the topology of features. Subsequently, PFR [26] employs a pseudo-set training strategy with frequency-based feature refinement to mitigate the lack of discriminative feature learning. Also, some models like FACT [5] try to use virtual prototypes to reserve for new ones, thus ensuring incremental learning ability. The SOTA method BiDist [7] utilizes a novel distillation structure to alleviate the effects of forgetting.
Recently, with the breakthroughs in 2D PTMs, a number of works have attempted to leverage the vast knowledge acquired by 2D PTMs, which is effective in learning new concepts and alleviating the problem of forgetting [11, 16]. With almost perfect performance, these models have attracted a lot of interest and attention.
The preceding methods are all 2D methods, except for Chowdhury et al. [8]’s work which explores the FSCIL task in 3D point cloud data. Chowdhury et al. [8] points out the domain gap between synthetic instances from the base task and real-scanned samples in incremental steps is a specific issue of 3D FSCIL. The domain gap exacerbates the inherent issues of catastrophic forgetting and overfitting in the FSCIL task. To address this, they propose microshapes to minimize variation between synthetic and real data and achieve domain adaptation.
However, Chowdhury et al. [8]’s work neither supplements knowledge for few-shot data and domain gap nor addresses the high-noise nature of real-world scanned data, resulting in limited performance in 3D FSCIL tasks. In contrast, our FILP-3D successfully addresses the shortcomings mentioned above by introducing V-L PTMs and two newly proposed modules.
3 Notations Table
In this section we first provide the notation of symbols involved in Sec. 4 and Sec. 5, as shown in Tab. 1. Note that in Sec. 4, $\mathcal{D}$ refers to the training datasets, whereas in Sec. 5, $\mathcal{D}$ refers to the testing datasets. And in Sec. 4 (except Section 4.4.1 Details of Pre-Processing), we replace $\mathbf{x}_{i}^{b}$, $y_{i}^{b}$ with $\mathbf{x}$, $y$ for clearer explanation.
Table 1: Notations table.
Symbols in Dataset
$B$ the number of task
$P$ the number of points contained in a point cloud
$\mathcal{D}$ the set of dataset in each session
$\mathcal{D}^{b}$ the dataset of the $b$-th session
$n_{b}$ the number of point cloud in $\mathcal{D}^{b}$
$\mathbf{x}_{i}^{b}$ the $i$-th point cloud in $\mathcal{D}^{b}$
$y_{i}^{b}$ the label of $i$-th point cloud in $\mathcal{D}^{b}$
$Y_{b}$ the label set of the $b$-th session
$K_{b}$ the number of visible classes for the $b$-th task
Symbols in Method
$N$ the number of views in the multi-view render stage
$C$ the embedding dimension of V-L PTM
$M$ the number of principal components
$K$ the number of the visual classes in this session
$\mathbf{f}^{p}$ the raw 3D feature obtained from the input point cloud $\mathbf{x}$ via the 3D encoder
$\mathbf{f}^{3D}$ the 3D feature modified from the raw 3D features $\mathbf{f}^{3D}$ via the adapter
${\rm\mathbf{D}}$ the multi-view depth maps rendered from the input point cloud $\mathbf{x}$
${\rm\mathbf{F}}^{d}$ the multi-view depth features obtained from the multi-view depth maps ${\rm\mathbf{D}}$ via the V-L PTM
$\mathbf{f}^{2D}$ the global 2D feature obtained from the multi-view depth maps ${\rm\mathbf{F}}^{d}$ via the Merger
$t_{k}$ the label name of $k$-th class
${\rm\mathbf{F}}^{t}$ the raw text features obtained from visible label set via V-L PTM and the template text prompt:
$\mathbf{f}^{g}$ the global feature fused from the 3D features $\mathbf{f}^{3D}$ and the global 2D feature $\mathbf{f}^{2D}$ via the fusion fuction
$\mathbf{V}$ the pre-processed principal components used to eliminate redundant dimensions
$\widetilde{\mathbf{f}}^{g}$ the global feature with redundant dimensions eliminated
$\widetilde{\mathbf{F}}^{t}$ the text features with redundant dimensions eliminated
$l_{k}$ the logit for class $k$-th class
$\mathbf{p}$ the ultimate prediction probability
$N_{Aug}$ the number of augmentation iterations during the contrastive learning stage
Symbols in Benchmark
${\rm Acc_{i}}$ the accuracy of $i$-th session.
${\Delta}$ relative accuracy dropping rate, where ${\rm\Delta=\frac{|Acc_{B}-Acc_{1}|}{Acc_{1}}}$
$Acc_{i}$ the accuracy of the $i$-th class
@@ 4.1 Problem Formulation
Assuming a sequence of $B$ tasks $\mathcal{D}=\{\mathcal{D}^{1},\mathcal{D}^{2},\dots,\mathcal{D}^{B}\}$, FSCIL methods incrementally recognize novel classes with a small amount of training data. For the $b$-th task $\mathcal{D}^{b}=\{(\mathbf{x}_{i}^{b},y_{i}^{b})\}_{i=1}^{n_{b}}$, we have $n_{b}$ training samples, in which an instance $\mathbf{x}_{i}^{b}$ has a class label $y_{i}^{b}$ and $y_{i}^{b}$ belongs to the label set $Y_{b}$. We stipulate that $Y_{b}\cap Y_{b^{\prime}}=\emptyset$ when $b\not=b^{\prime}$. After the $b$-th training task, trained models are required to classify test sets of all the previous tasks $\{\mathcal{D}^{1},\dots,\mathcal{D}^{b}\}$. Note that, in the 3D FSCIL setting, $\mathbf{x}_{i}^{b}\in\mathbb{R}^{P\times 3}$ denotes a point cloud object, where $P$ denotes the number of points in a point cloud object. $\mathcal{D}^{1}$ is the base task with a large-scale 3D training dataset, while much fewer samples are included in the incremental task $\mathcal{D}^{b}$, i.e., $n_{b}\ll n_{1}$ for $b>1$.
@@ 4.2 Model Overview
Figure 2: Overview of FILP-3D. FILP-3D mainly consists of three components, i.e., 3D branch (SNC), CLIP backbone, and the classify component (RFE). The SNC generates 3D feature $\mathbf{f}^{3D}$ from the input point cloud. The CLIP backbone also generates global 2D feature $\mathbf{f}^{2D}$ and text features ${\rm\mathbf{F}}^{t}$ from the input point cloud and the class names respectively. Then, the 3D feature $\mathbf{f}^{3D}$ and the global 2D feature $\mathbf{f}^{2D}$ will be fused as a global feature $\mathbf{f}^{g}$, and be used to calculate probability alongside text features after redundant dimensions eliminated in the RFE module.
The framework of FILP-3D is shown in Figure 2. For each training sample $(\mathbf{x},y)$, we mainly have three branches: 1) point cloud $\mathbf{x}$ is fed into 3D encoder to obtain an original point feature $\mathbf{f}^{p}$ and then adapted to the 3D feature $\mathbf{f}^{3D}$, 2) $\mathbf{x}$ is then rendered as multi-view depth maps ${\rm\mathbf{D}}$, embedded as multi-view depth features ${\rm\mathbf{F}}^{d}$, and finally merged into a global 2D feature $\mathbf{f}^{2D}$, 3) visible class names are templated into text prompts and encoded as text features ${\rm\mathbf{F}}^{t}$. The 3D features $\mathbf{f}^{3D}$ and the global 2D feature $\mathbf{f}^{2D}$ are then fused as a global feature $\mathbf{f}^{g}$. Afterward, we use pre-processed principal components $\mathbf{V}$ (details of pre-processing will be mentioned in Sec. 4.4.1) to eliminate redundant dimensions of $\mathbf{f}^{g}$ and $\mathbf{F}^{t}$ and generate $\widetilde{\mathbf{f}}^{g}$ and $\widetilde{\mathbf{F}}^{t}$. The final logits $l$ are calculated by our proposed renormalized cosine similarity and the predicted probability $\mathbf{p}$ is the result of applying the softmax function to the logits $l$.